OpenStudy (anonymous):

find the 20th term of {270,90,30,10,....] I found the ratio is 3 what do I do find the 20th term

OpenStudy (anonymous):

\[270\times(\frac{1}{3})^{19}\]

OpenStudy (anonymous):

or if you prefer \[\frac{270}{3^{19}}\] btw the common ratio is \(\frac{1}{3}\) and not \(3\)

OpenStudy (anonymous):

can you draw it out I cant understand what you typed

OpenStudy (anonymous):

nevermind got it

OpenStudy (anonymous):

thank you